# Causal-Adapter: Taming Text-to-Image Diffusion for Faithful Counterfactual Generation

### v3

@@ 3 Experiments
We conduct extensive experiments on three counterfactual image generation datasets across different domains: the synthetic Pendulum dataset (Yang et al., 2021), the human-face dataset CelebA (Liu et al., 2015), and the medical imaging dataset Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Petersen et al., 2010). We follow the benchmarking experimental settings (Melistas et al., 2024; Komanduri et al., 2024c) for fair comparison. During evaluation, we follow the official setups for each of the datasets and report: (1) Effectiveness: whether the intervention succeeds, measured by pretrained classifiers using F1-score or MAE depending on the attribute type. (2) Composition: reconstruction quality under null interventions, measured by MAE and LPIPS distance. (3) Realism: visual quality of counterfactual images, evaluated with Fréchet Inception Distance (FID). (4) Minimality: non-intervened attributes remain minimally affected, assessed with the Counterfactual Latent Divergence (CLD) metric. Implementation details are presented in Appendix C.
@@ 3.1 Results
Synthetic Imaging Counterfactuals.
Figure 5: Pendulum counterfactuals with traversal editing along each attribute.
We first evaluate our approach on the Pendulum dataset, which consists of four continuous causal variables (pendulum angle, light position, shadow length, and shadow position). We compare against CausalVAE (Yang et al., 2021), as well as diffusion-based methods DisDiffAE (Preechakul et al., 2022) and CausalDiffAE (Komanduri et al., 2024c), using the causal graph shown in Figure 5. As reported in Table 1, our method (row 4) achieves state-of-the-art intervention performance across most attributes. In particular, under $do(l)$ (light intervention), we obtain up to a 91% reduction in MAE for light prediction (from 0.434 to 0.041), indicating highly accurate control over light movement. Moreover, when intervening on light, our approach correctly preserves the pendulum angle while inducing causal changes in the descendant shadow attributes (shadow length and position), consistent with the real-world physical law. We further conduct an investigation into the role of causal mechanisms. Without explicit causal mechanisms (row 5), the model exhibits larger intervention errors, underscoring the necessity of updating semantic variables with causal dependencies during generation. In contrast, injecting synthetic ground-truth labels (row 6) yields performance close to ours (row 4), indicating that our causal mechanism injection provides a principled approximation to ground-truth causal reasoning, generating precise and faithful counterfactuals over all attributes as shown in Figure 5.
Table 1: Intervention effectiveness on Pendulum test set. We report MAE from pretrained regressors under four interventions. w/ CM: with causal mechanisms; w/o CM: without causal mechanisms; w/ GT: ground-truth labels injected. $\sim$” denotes descendant attributes remain unaffected.
@@
1. CausalVAE 24.86 23.03 20.47 11.58 34.20 26.01 35.49 47.06 1.946 1.430 2.020 1.720 52.52 72.50 57.03 32.78
2. DisDiffAE 0.668 0.648 0.647 0.647 0.656 0.654 0.630 0.651 0.550 0.527 0.560 0.516 0.474 0.475 0.479 0.534
3. CausalDiffAE 0.297 0.132 0.031 0.034 0.045 0.434 0.035 0.064 0.136 0.322 0.492 0.082 0.146 0.303 0.064 0.471
4. $\text{Ours}_{\text{w/~CM}}$ 0.014 0.035 0.043 0.259 0.045 0.041 0.058 0.120 0.028 0.051 0.489 0.110 0.030 0.033 0.041 0.336
5. $\text{Ours}_{\text{w/o~CM}}$ 0.159 0.183 $\sim$ $\sim$ 0.060 0.173 $\sim$ $\sim$ 0.143 0.235 $\sim$ $\sim$ 0.086 0.155 $\sim$ $\sim$
6. $\text{Ours}_{\text{w/~GT}}$ 0.013 0.033 $\sim$ $\sim$ 0.043 0.039 $\sim$ $\sim$ 0.025 0.036 $\sim$ $\sim$ 0.028 0.035 $\sim$ $\sim$
Table 2: Intervention effectiveness on CelebA test set. F1 scores under four interventions.
1. VAE 35.0 78.2 81.6 81.9 97.7 90.9 95.9 97.3 94.4 82.8 29.6 94.5 2.3 49.6 4.5 41.2
2. HVAE 65.4 89.3 90.8 89.9 98.8 94.9 99.4 95.0 95.2 95.1 44.1 91.6 2.0 86.0 4.5 61.1
3. GAN 41.3 71.0 81.8 79.9 95.2 98.2 92.0 96.1 90.8 83.8 23.3 90.7 2.1 82.0 5.5 49.2
4. Ours 58.4 89.6 93.9 89.4 99.6 99.9 74.4 92.3 99.7 96.1 51.9 98.1 58.8 91.0 76.4 58.5
@@
5. Ours 57.2 88.2 94.5 81.9 99.6 99.7 73.4 89.1 99.6 96.8 48.3 96.5 37.3 78.0 46.4 51.7
Figure 6: Composition, realism and minimality results on CelebA test set. “-” indicates attention guidance was not applied for reconstruction.
Method Composition Realism Minimality
MAE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ CLD $\downarrow$
1. VAE 18.695 0.282 59.397 0.299
2. HVAE 7.143 0.122 35.712 0.305
3. GAN 60.120 0.276 27.861 0.304
4. Ours 3.535 0.017 8.509 0.309
with attention guidance for better realism and minimality
5. Ours - - 5.386 0.301
Figure 7: CelebA counterfactuals from Causal-Adapter compared with prior methods.
@@
Following the benchmarking of Melistas et al. (2024), we evaluate Causal-Adapter on CelebA test set for human face counterfactual generation across four categorical attributes (age, gender, beard, bald) with the causal graph shown in Figure 1. We also incorporate attention guidance to perform localized editing and assess the utility of our learned attention maps. Table 2 reports intervention effectiveness, while Table 7 summarizes composition, realism, and minimality. Compared with VAE, HVAE, and GAN, our method achieves best performance across most interventions, including up to an $86\%$ reduction in LPIPS (composition, from 0.122 to 0.017) and an $58\%$ reduction in FID (realism, from 27.861 to 8.509). HVAE achieves competitive effectiveness via post-training classifier optimization but produces classifier-biased artifacts and reduced fidelity (Figure 7). Besides, our method delivers significant improvements under $do(a)$ and $do(br)$ on F1 score of bald attribute, demonstrating causal faithfulness in representing baldness across diverse individuals. As illustrated in Figure 7, Causal-Adapter can successfully add baldness to both male and female under $do(bl)$, and under interventions on causal parents such as $do(a)$, it can jointly edit age and baldness, whereas baselines fail to maintain causal consistency. Further localized editing with attention guidance balances intervention effectiveness and identity preservation, achieving the best FID (from 8.509 to 5.386). Qualitative results further confirm that the learned attention maps enable precise, localized detailed editing (e.g., modifying gender without altering hairstyle), enable Causal-Adapter to preserve core identity while enforcing causal interventions.
@@
1. VAE 0.170.03 0.150.06 0.150.06 0.080.05 0.200.04 0.080.05 0.520.15 0.480.15 0.460.10 18.88 0.306 278.245 0.352
2. HVAE 0.090.03 0.120.06 0.130.06 0.060.04 0.040.01 0.060.04 0.380.15 0.410.16 0.410.11 3.38 0.101 74.696 0.347
3. GAN 0.170.02 0.160.07 0.160.06 0.120.02 0.220.03 0.120.03 0.140.03 0.160.03 0.050.02 24.26 0.268 113.749 0.353
4. Ours 0.090.01 0.110.03 0.110.03 0.030.01 0.030.01 0.030.01 0.530.09 0.550.09 0.480.06 3.54 0.035 9.130 0.346
@@
Brain Imaging Counterfactuals
We further evaluate our method on ADNI dataset, which includes six attributes in both categorical variables (ApoE, Sex, Slice) and continuous variables (Age, Brain Volume, Ventricular Volume). Following Melistas et al. (2024), we intervene on three generative conditioning attributes (Brain Volume, Ventricular Volume, Slice) and report the results in Table 3. Our approach achieves best performance in intervention effectiveness (up to $50\%$ MAE reduction in fine-grained edits) and minimality, while also delivering strong realism ($87\%$ FID reduction) even without attention guidance. Qualitative results in Figure 8 further show that our model produces sharp and localized interventional changes consistent with the causal graph. For example, edits to ApoE, Age, or Sex appropriately influence Brain and Ventricular Volumes. In particular, fine-grained edits to Ventricular Volume visibly enlarge the ventricle region while faithfully preserving subject identity.
@@
We conduct an ablation study on the CelebA validation set to evaluate the contribution of each regularizer in Causal-Adapter. As shown in Figures 9a and b, the plain adapter achieves the lowest intervention effectiveness. Adding the PAI module yields a consistent average gain of $11.6\%$ F1 across all attributes with slight increases in FID and CLD, indicating that aligning causal semantics with spatial features in the diffusion latents enables more effective edits. Incorporating CTC further improves intervention effectiveness by enforcing token embedding disentanglement, while also reducing CLD (from 0.317 to 0.310) and keeping FID stable (from 8.453 to 8.643), resulting more faithful counterfactual generation. Figure 9c presents qualitative ablation results. Spurious correlations arise in plain adapter, e.g., beard interventions in females induce male facial features (c2). PAI alleviates attribute entanglement (c3), while the complete regularization with CTC produces bearded female, demonstrating mitigation of such correlations (c4). Under bald interventions, the baselines alter skin color or age (c2–c3), whereas the full regularization edits baldness with only minor facial changes (c4). Finally, attention guidance enhances identity preservation with localizing edits (c5). Extended results are presented in Appendix D.
Figure 9: Ablation study on CelebA validation set. (a) Average intervention effectiveness. (b) Realism and minimality. (c) Qualitative examples, with dotted boxes indicating results of localized editing.
@@ 4 Conclusion
We introduced Causal-Adapter to tame Text-to-Image diffusion models for counterfactual image generation. Our motivational study revealed that current Text-to-Image diffusion model based editing approaches lack an explicit structural causal model for attribute control and rely heavily on prompt engineering, making it difficult to generate faithful counterfactual images. In contrast, Causal-Adapter is a simple yet effective framework that leverages a frozen diffusion backbone and injects causal semantic attributes through a pluggable adapter network to explicitly learn causal semantics. We further proposed prompt aligned injection and conditioned token contrastive optimization, which align attribute semantics with spatial features and promote disentanglement in the latent space, reducing spurious correlations while preserving identity for generations. Extensive experiments across different tasks and datasets demonstrate our approach’s superior performance, confirming its robustness and adaptability.
Ethics Statement. This work makes use of two real-world datasets: CelebA (Liu et al., 2015) and ADNI (Petersen et al., 2010), together with pretrained text-to-image generation models. We emphasize that our research is conducted strictly for scientific purposes, and we strongly condemn any misuse of generative AI to produce content that harms individuals, violates privacy, or spreads misinformation. While our approach demonstrates capabilities in generating human faces and MRI images, we acknowledge the potential for misuse. To mitigate these risks, we uphold the highest ethical standards, including adherence to applicable legal and institutional frameworks, respect for data privacy, and a commitment to promoting socially beneficial applications of generative models.
Reproducibility Statement. In Section 2, we provide detailed formulations, illustrative examples, and visual demonstrations (e.g., Figure 4) to clarify the model structure and mechanisms. Section 3 and Appendix C describe the training datasets, model parameters, and implementation components. Comprehensive quantitative and qualitative results are reported across multiple domains, including Table 1, Table 2, Figure 7, and Figure 8, with comparisons to existing baselines under standard benchmarks. All experiments are conducted on publicly available datasets (CelebA and ADNI) and evaluated with widely used metrics such as FID and LPIPS. To further promote reproducibility, we will release the experimental code upon acceptance of this work.
Appendix A Extended Related Works
@@
aims to synthesize images that reflect the visual effect of a hypothetical intervention while preserving non-intervened attributes according to an underlying causal graph and maintaining instance-specific identity details (Komanduri et al., 2024a; Melistas et al., 2024). Existing approaches often augment generative models with explicit structural causal models (SCMs) and follow Pearl’s abduction–action–prediction paradigm (Pearl, 2009; 2013; Pawlowski et al., 2020; Shen et al., 2022; Sanchez & Tsaftaris, 2022; De Sousa Ribeiro et al., 2023; Wu et al., 2025; Spyrou et al., 2025). Early work predominantly relied on VAEs or GANs (Kingma & Welling, 2013; Goodfellow et al., 2014; Kocaoglu et al., 2018; Pawlowski et al., 2020), as their noise-injection and variational sampling mechanisms offered a natural way to represent exogenous uncertainty, while their objectives encouraged disentanglement of causal factors, inspired by $\beta$-VAE (Higgins et al., 2017). De Sousa Ribeiro et al. (2023) extends this line by incorporating hierarchical VAEs (Vahdat & Kautz, 2020) to estimate direct, indirect, and total causal effects, whereas Wu et al. (2025) combined variational Bayesian inference with adversarial training to improve abduction and preserve identity. Other works (Yang et al., 2021; Shen et al., 2022; Komanduri et al., 2024b; Li et al., 2024) integrates SCM priors directly into the latent space of VAEs, enabling implicit causal reasoning on image encodings. However, variational optimization inevitably introduces uncertainty into the learned representations, which can lead to posterior collapse or the loss of semantically meaningful factors. This results in an inherent trade-off between high-fidelity image synthesis and flexible attribute control (Higgins et al., 2017; Alemi et al., 2018). Sanchez & Tsaftaris (2022) introduces DiffSCM, the first framework to integrate diffusion with SCMs for counterfactual generation. DiffSCM employed DDIM inversion (Song et al., 2021) for abduction and conditioned the generative process on causal labels, but it was limited to small parent sets and simple causal structures. Pan & Bareinboim (2024) extends DiffSCM by combining a VAE for preliminary counterfactual generation with diffusion-based refinement, shifting the conditioning from attribute labels to pre-generated images to produce refined counterfactuals. Recent works (Komanduri et al., 2024c; Rasal et al., 2025; Xia et al., 2025) build on Diffusion Autoencoder framework (Preechakul et al., 2022), equipping diffusion models with variational encoders to inject semantic attributes into diffusion latents. However, these methods require heavy post-training or fine-tuning and remain prone to spurious correlations. A key limitation is that disentanglement is typically enforced through auxiliary encoders with limited influence on the intrinsic latent space of diffusion models, failing to align semantic attributes with disentangled spatial representations of images and leaving causal disentanglement incomplete (Wu et al., 2025; Yang et al., 2024). To address these limitations, we propose Causal-Adapter, an adaptive and modular framework that employs an adapter encoder to explicitly learn causal interactions between semantic attributes. We further introduce two regularization strategies: Prompt Aligned Injection (PAI) and Conditioned Token Contrastive Loss (CTC). These strategies align semantic attributes with spatial features in the diffusion latents and separate token embeddings across conditions, thereby enhancing causal representation learning and reducing spurious correlations, all while preserving the pre-trained diffusion weights.
@@
aims to manipulate existing images according to user-provided natural language instructions (Brooks et al., 2023). Most approaches rely on an inversion process (Song et al., 2021), where the image is projected into a latent state and then synthesized under modified conditions (Hertz et al., 2023; Ho & Salimans, 2021). However, DDIM inversion and classifier-free guidance often often interfere with each other, leading to a trade-off between preserving essential content and achieving faithful edits (Ju et al., 2024; Huberman-Spiegelglas et al., 2024; Kynkäänniemi et al., 2024). To mitigate this, Null-text inversion (Mokady et al., 2023) learns a null embedding to account for reconstruction discrepancies, while  Dong et al. (2023) optimizes conditional embeddings to reduce information loss in unconditional guidance. Yet, both methods require costly per-sample optimization. Subsequent works (Ju et al., 2024; Xu et al., 2024; Miyake et al., 2025) improve efficiency by recording residual losses between conditional and unconditional embeddings and re-injecting them during editing, to stabilize edits and preserving content. Textual inversion (Gal et al., 2023) improves content preservation by disentangling single concepts: it learns new text embeddings from a few personalized images to represent objects in novel contexts. Extending this, Vinker et al. (2023) decomposes concept embeddings into sub-concepts via learned vectors injected into the latent space of T2I model, while Jin et al. (2024) performs multi-concept disentanglement using adjective-based prompts and contrastive optimization. Despite these advances, generic T2I based editing remains inadequate for counterfactual generation. Current methods rely heavily on prompt engineering to achieve reliable edits and lack explicit, learnable SCM over semantic attributes. As a result, they struggle to ensure both causal faithfulness and identity preservation in counterfactual image generation.

### v5

@@ 3 Experiments
We conduct experiments on four counterfactual image generation datasets across different domains: the synthetic Pendulum dataset (Yang et al., 2021), the human-face dataset CelebA (Liu et al., 2015) and its high-resolution restoration CelebA-HQ (Karras et al., 2017), and the medical imaging dataset Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Petersen et al., 2010). We follow the benchmarking experimental settings (Melistas et al., 2024; Komanduri et al., 2024c; Rasal et al., 2025) for fair comparison. During evaluation, we follow the official setups for each of the datasets and report: (1) Effectiveness: whether the intervention succeeds, measured by pretrained classifiers using F1-score or MAE depending on the attribute type. (2) Composition: reconstruction quality under null interventions, measured by MAE and LPIPS distance. (3) Realism: visual quality of counterfactual images, evaluated with Fréchet Inception Distance (FID). (4) Minimality: non-intervened attributes remain minimally affected, assessed with the Counterfactual Latent Divergence (CLD) metric. Implementation details are presented in Appendix C.
Figure 5: Qualitative results on the Pendulum dataset demonstrating counterfactual traversals. When intervening on the pendulum angle or light position, our approach induces causally consistent changes in the descendant shadow attributes (length and position), in accordance with the underlying causal laws.
@@ 3.1 Results
Table 1: Intervention effectiveness on Pendulum test set. We report MAE from pretrained regressors under four interventions. w/ CM: with causal mechanisms; w/o CM: without causal mechanisms; w/ GT: ground-truth labels injected. $\sim$” denotes descendant attributes remain unaffected. The table follows the causal graph: e.g., under the “Pendulum $(p)$ MAE”, $do(p)$ or $do(l)$ report the pendulum MAE after intervening on $p$ or $l$.
@@
2 CausalVAE 24.86 23.03 20.47 11.58 34.20 26.01 35.49 47.06 1.946 1.430 2.020 1.720 52.52 72.50 57.03 32.78
3 DisDiffAE 0.668 0.648 0.647 0.647 0.656 0.654 0.630 0.651 0.550 0.527 0.560 0.516 0.474 0.475 0.479 0.534
4 CausalDiffAE 0.297 0.132 0.031 0.034 0.045 0.434 0.035 0.064 0.136 0.322 0.492 0.082 0.146 0.303 0.064 0.471
5 $\text{Ours}_{\text{w/~CM}}$ 0.014 0.035 0.043 0.259 0.045 0.041 0.058 0.120 0.028 0.051 0.489 0.110 0.030 0.033 0.041 0.336
6 $\text{Ours}_{\text{w/o~CM}}$ 0.159 0.183 $\sim$ $\sim$ 0.060 0.173 $\sim$ $\sim$ 0.143 0.235 $\sim$ $\sim$ 0.086 0.155 $\sim$ $\sim$
7 $\text{Ours}_{\text{w/~GT}}$ 0.013 0.033 $\sim$ $\sim$ 0.043 0.039 $\sim$ $\sim$ 0.025 0.036 $\sim$ $\sim$ 0.028 0.035 $\sim$ $\sim$
Table 2: CelebA results. We report target attribute effectiveness and include composition, realism, and minimality metrics.
[…]
$do(a)$ $do(g)$ $do(br)$ $do(bl)$ (Comp.) (Real.) (Min.)
2 VAE 35.0 90.9 29.6 41.2 0.282 59.397 0.299
3 HVAE 65.4 94.9 44.1 61.1 0.122 35.712 0.305
4 GAN 41.3 98.2 23.3 49.2 0.276 27.861 0.304
5 Ours 58.5 99.9 52.1 58.8 0.017 8.152 0.310
@@
7 Ours (AG) 57.1 99.7 48.2 51.4 5.213 0.301
Synthetic Imaging Counterfactuals.
We first evaluate our approach on the Pendulum dataset, which consists of four continuous causal variables (pendulum angle, light position, shadow length, and shadow position). We compare against CausalVAE (Yang et al., 2021), as well as diffusion-based methods DisDiffAE (Preechakul et al., 2022) and CausalDiffAE (Komanduri et al., 2024c), using the causal graph shown in Figure 5. As reported in Table 1, our method (row 4) achieves state-of-the-art intervention performance across most attributes. In particular, under $do(l)$ (light intervention), we obtain up to a 91% reduction in MAE for light prediction (from 0.434 to 0.041), indicating highly accurate control over light movement. Moreover, when intervening on light, our approach correctly preserves the pendulum angle while inducing causal changes in the descendant shadow attributes (shadow length and position), consistent with the real-world physical law. We further conduct an investigation into the role of causal mechanisms. Without explicit causal mechanisms (row 5), the model exhibits larger intervention errors, underscoring the necessity of updating semantic variables with causal dependencies during generation. In contrast, injecting synthetic ground-truth labels (row 6) yields performance close to ours (row 4), indicating that our causal mechanism provides a principled approximation to ground-truth causal reasoning, generating faithful counterfactuals over all attributes as shown in Figure 5.
@@
Following the benchmarking of Melistas et al. (2024), we evaluate Causal-Adapter on CelebA test set for human face counterfactual generation across four categorical attributes (age, gender, beard, bald) with the causal graph shown in Figure 1. We also incorporate attention guidance to perform localized editing and assess the utility of our learned attention maps. Table 2 reports intervention effectiveness, composition, realism, and minimality. Our method achieves best performance across most interventions, including up to an $86\%$ reduction in LPIPS (composition, from 0.122 to 0.017) and an $79\%$ reduction in FID (realism, from 27.861 to 8.152). HVAE achieves competitive effectiveness via post-training classifier optimization but produces classifier-biased artifacts and reduced fidelity (Figure 6). Further localized editing with attention guidance balances intervention effectiveness and identity preservation, achieving the best FID (from 8.152 to 5.213). Qualitative results confirm that the learned attention maps enable precise, localized detailed editing (e.g., modifying gender without altering hairstyle), enable Causal-Adapter to preserve core identity while enforcing causal interventions. Full results are in Appendix E.2.
Figure 6: Qualitative comparison on the CelebA dataset. The results demonstrate that our method preserves core identity features while effectively enforcing causal interventions.
Table 3: Intervention target effectiveness on ADNI.
@@
$do(b)$ $do(v)$ $do(s)$ (Comp.) (Real.) (Min.)
2 VAE 0.17 0.20 0.46 0.306 278.245 0.352
3 HVAE 0.09 0.04 0.41 0.101 74.696 0.347
4 GAN 0.17 0.22 0.05 0.268 113.749 0.353
5 Ours 0.09 0.03 0.48 0.035 9.130 0.346
@@
Table 4: CelebA-HQ results. Intervention effectiveness, reversibility, and identity preservation for eyeglasses and smiling.
[…]
$do(g)$ (Rev.) (IDP) $do(s)$ (Rev.) (IDP)
2 VCI 3.39 33.81
3 HVAE 65.31 75.33
4 DiffCounter 96.86 0.185 0.096 94.93 0.183 0.066
5 Ours 99.26 0.049 0.084 94.15 0.028 0.028
6 Ours (DiT) 97.39 0.086 0.060 94.71 0.089 0.035
Brain Imaging Counterfactuals.
We further evaluate our method on ADNI dataset, which includes six attributes in both categorical variables (ApoE, Sex, Slice) and continuous variables (Age, Brain Volume, Ventricular Volume). Following Melistas et al. (2024), we intervene on three generative conditioning attributes (Brain Volume, Ventricular Volume, Slice) and report the results in Table 3. Our approach achieves best performance in intervention effectiveness (up to $50\%$ MAE reduction in fine-grained edits) and minimality, while also delivering strong realism ($87\%$ FID reduction) even without attention guidance. Qualitative results (Figure 7) further show that our model produces sharp and localized interventional changes consistent with the causal graph. In particular, fine-grained edits to Ventricular Volume visibly enlarge the ventricle region while preserving subject identity. Full results are in Appendix E.3.
High-Resolution Face Counterfactuals.
We further compare Causal-Adapter with recent SOTA methods, VCI (Wu et al., 2025), HVAE (De Sousa Ribeiro et al., 2023), and DiffCounter (Rasal et al., 2025) on high-resolution human face dataset CelebA-HQ. We follow the same settings in DiffCounter for fair comparison, focusing on three categorical variables (eyeglass, smiling, mouth-open). Quantitative results are reported in Table 4, with qualitative examples shown in Figure 8. Our method (row 4) achieves intervention effectiveness comparable to DiffCounter, while substantially improving both reversibility and identity preservation. We further evaluate generalization by instantiating Causal-Adapter with a diffusion transformer (DiT) backbone (Esser et al., 2024). As shown in Table 4 (row 5) and Appendix E.4, Causal-Adapter transfers effectively across distinct T2I backbones and produces causally faithful, high-resolution counterfactuals.
Figure 8: CelebA-HQ counterfactuals, together with reversible cases that map back to the original observations, illustrating identity preservation and reversibility.
Figure 9: Ablation study on CelebA validation set. (a) Average intervention effectiveness. (b) Realism and minimality. (c) Qualitative examples, with dotted boxes indicating results of localized editing. Both quantitative and qualitative results demonstrate the necessity of the proposed PAI and CTC.
@@
We conduct an ablation study on the CelebA validation set to evaluate the contribution of each regularizer. As shown in Figures 9(a) and (b), the plain adapter achieves the lowest intervention effectiveness. Adding the PAI module yields a consistent average gain of $11.6\%$ F1 across all attributes with slight increases in FID and CLD, indicating that aligning causal semantics with spatial features in the diffusion latents enables more effective edits. Incorporating CTC further improves intervention effectiveness by enforcing token embedding disentanglement, while also reducing CLD (from 0.317 to 0.310) and keeping FID stable (from 8.453 to 8.643), resulting more faithful counterfactual generation. Figure 9c presents qualitative ablation results. Spurious correlations arise in plain adapter, e.g., beard interventions in females induce male facial features (c2). PAI alleviates attribute entanglement (c3), while the complete regularization with CTC produces bearded female, demonstrating mitigation of such correlations (c4). Under bald interventions, the baselines alter skin color or age (c2–c3), whereas the full regularization edits baldness with only minor facial changes (c4). Finally, attention guidance enhances identity preservation with localizing edits (c5). Extended results are presented in Appendix D.
Beyond Pre-defined Graphs.
We additionally explore incorporating causal structure learning (Appendix G) as a preliminary attempt to learn the causal graph from scratch.
@@ 4 Conclusion
We introduced Causal-Adapter, a modular framework to tame Text-to-Image diffusion models for counterfactual generation. Our motivational study revealed that current Text-to-Image diffusion based editing approaches lack an explicit structural causal model for attribute control, making it difficult to generate faithful counterfactual images. In contrast, Causal-Adapter is a simple yet effective framework that leverages a frozen diffusion backbone and injects causal semantic attributes through a pluggable adapter network to explicitly learn causal semantics. We further proposed prompt aligned injection and conditioned token contrastive optimization, which align attribute semantics with spatial features and promote disentanglement in the latent space, reducing spurious correlations while preserving identity for generations. Causal-adapter achieves superior counterfactual generation performance on multiple datasets. Extensive evaluation across diverse settings further confirm that Causal-Adapter provides a robust, scalable, and practical alternative for enabling causal editing in modern T2I systems.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix Table of Contents
  • Appendix A - Extended Related Works ........................................................................................................................................................................A
  • Appendix B - Full Motivational Study Results ........................................................................................................................................................................B
  • Appendix C - Dataset and Implementation ........................................................................................................................................................................C
    • C.1: Implementation Details ........................................................................................................................................................................C.1
    • C.2: Metrics ........................................................................................................................................................................C.2
    • C.3: Full Regularization Algorithm ........................................................................................................................................................................C.3
  • Appendix D - Extended Ablation Results ........................................................................................................................................................................D
    • D.1: Effect of Guidance Scale ........................................................................................................................................................................D.1
    • D.2: Effect of DDIM Steps ........................................................................................................................................................................D.2
    • D.3: Investigation of Attention Guidance ........................................................................................................................................................................D.3
    • D.4: Qualitative Evidence ........................................................................................................................................................................D.4
  • Appendix E - Additional Results ........................................................................................................................................................................E
    • E.1: Pendulum ........................................................................................................................................................................E.1
    • E.2: CelebA ........................................................................................................................................................................E.2
    • E.3: ADNI ........................................................................................................................................................................E.3
    • E.4: CelebA-HQ ........................................................................................................................................................................E.4
    • E.5: Attention Maps ........................................................................................................................................................................E.5
  • Appendix F - Counterfactual Identifiability ........................................................................................................................................................................F
  • Appendix G - Assumption of a Known Causal Graph ........................................................................................................................................................................G
Appendix A Extended Related Works [Back to Contents]
@@
aims to synthesize images that reflect the visual effect of a hypothetical intervention while preserving non-intervened attributes according to an underlying causal graph and maintaining instance-specific identity details (Komanduri et al., 2024a; Melistas et al., 2024). Existing approaches often augment generative models with explicit structural causal models (SCMs) and follow Pearl’s abduction–action–prediction paradigm (Pearl, 2009, 2013; Pawlowski et al., 2020; Shen et al., 2022; Sanchez and Tsaftaris, 2022; De Sousa Ribeiro et al., 2023; Wu et al., 2025; Spyrou et al., 2025). Early work predominantly relied on VAEs or GANs (Kingma and Welling, 2013; Goodfellow et al., 2014; Kocaoglu et al., 2018; Pawlowski et al., 2020), as their noise-injection and variational sampling mechanisms offered a natural way to represent exogenous uncertainty, while their objectives encouraged disentanglement of causal factors, inspired by $\beta$-VAE (Higgins et al., 2017). De Sousa Ribeiro et al. (2023) extends this line by incorporating hierarchical VAEs (Vahdat and Kautz, 2020) to estimate direct, indirect, and total causal effects, whereas Wu et al. (2025) combined variational Bayesian inference with adversarial training to improve abduction and preserve identity. Other works (Yang et al., 2021; Shen et al., 2022; Komanduri et al., 2024b; Li et al., 2024) integrates SCM priors directly into the latent space of VAEs, enabling implicit causal reasoning on image encodings. However, variational optimization inevitably introduces uncertainty into the learned representations, which can lead to posterior collapse or the loss of semantically meaningful factors. This results in an inherent trade-off between high-fidelity image synthesis and flexible attribute control (Higgins et al., 2017; Alemi et al., 2018). Sanchez and Tsaftaris (2022) introduces DiffSCM, the first framework to integrate diffusion with SCMs for counterfactual generation. DiffSCM employed DDIM inversion (Song et al., 2021) for abduction and conditioned the generative process on causal labels, but it was limited to small parent sets and simple causal structures. Pan and Bareinboim (2024) extends DiffSCM by combining a VAE for preliminary counterfactual generation with diffusion-based refinement, shifting the conditioning from attribute labels to pre-generated images to produce refined counterfactuals. Chao et al. (2024) models structural equations directly using diffusion processes, enabling counterfactual sampling on a predefined causal graph. While DCM also answers interventional queries, it trains a dedicated diffusion model per causal node. Recent works (Komanduri et al., 2024c; Rasal et al., 2025; Xia et al., 2025) build on Diffusion Autoencoder framework (Preechakul et al., 2022), equipping diffusion models with variational encoders to inject semantic attributes into diffusion latents. However, these methods require heavy post-training or fine-tuning and remain prone to spurious correlations. A key limitation is that disentanglement is typically enforced through auxiliary encoders with limited influence on the intrinsic latent space of diffusion models, failing to align semantic attributes with disentangled spatial representations of images and leaving causal disentanglement incomplete (Wu et al., 2025; Yang et al., 2024).
To address these limitations, we propose Causal-Adapter, an adaptive and modular framework that employs an adapter encoder to explicitly learn causal interactions between semantic attributes. We further introduce two regularization strategies: Prompt Aligned Injection (PAI) and Conditioned Token Contrastive Loss (CTC). These strategies align semantic attributes with spatial features in the diffusion latents and separate token embeddings across conditions, thereby enhancing causal representation learning and reducing spurious correlations, all while preserving the pre-trained diffusion weights.
@@
aims to manipulate existing images according to user-provided natural language instructions (Brooks et al., 2023). Most approaches rely on an inversion process (Song et al., 2021), where the image is projected into a latent state and then synthesized under modified conditions (Hertz et al., 2023; Ho and Salimans, 2021). However, DDIM inversion and classifier-free guidance often often interfere with each other, leading to a trade-off between preserving essential content and achieving faithful edits (Ju et al., 2024; Huberman-Spiegelglas et al., 2024; Kynkäänniemi et al., 2024). To mitigate this, Null-text inversion (Mokady et al., 2023) learns a null embedding to account for reconstruction discrepancies, while  Dong et al. (2023) optimizes conditional embeddings to reduce information loss in unconditional guidance. Yet, both methods require costly per-sample optimization. Subsequent works (Ju et al., 2024; Xu et al., 2024; Miyake et al., 2025) improve efficiency by recording residual losses between conditional and unconditional embeddings and re-injecting them during editing, to stabilize edits and preserving content. Textual inversion (Gal et al., 2023) improves content preservation by disentangling single concepts: it learns new text embeddings from a few personalized images to represent objects in novel contexts. Extending this, Vinker et al. (2023) decomposes concept embeddings into sub-concepts via learned vectors injected into the latent space of T2I model, while Jin et al. (2024) performs multi-concept disentanglement using adjective-based prompts and contrastive optimization. Lyu et al. (2024) propose a one-dimensional non-invasive adapter that modulates concept semi-permeability in frozen diffusion models for concept erosion, and Yeganeh et al. (2025) address domain shift in medical imaging, supporting counterfactual-like edits such as aging or disease progression without explicit structural constraints. Despite these advances, generic T2I editing remains insufficient for causal counterfactual generation. Existing methods depend heavily on prompt manipulation and do not incorporate a learnable structural causal model (SCM). They make no use of observed semantic attributes or a predefined causal graph, and therefore cannot enforce that edits follow correct causal dependencies. As a result, current T2I-based editing techniques struggle to simultaneously maintain causal faithfulness and identity preservation in counterfactual image generation.